### v1

# Optimal Control of the Future via Prospective Foraging

[…]
figurec
Optimal Control of the Future via Prospective Foraging
Yuxin Bai Johns Hopkins University Aranyak Acharyya Johns Hopkins University Ashwin De Silva Johns Hopkins University Zeyu Shen Johns Hopkins University James Hassett Johns Hopkins University Joshua T. Vogelstein Johns Hopkins University
Optimal control of the future is the next frontier for AI. Current approaches to this problem are typically rooted in either reinforcement learning or online learning. While powerful, these frameworks for learning are mathematically distinct from Probably Approximately Correct (PAC) learning, which has been the workhorse for the recent technological achievements in AI. We therefore build on the prior work of prospective learning, an extension of PAC learning (without control) in non-stationary environments [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik]. Here, we further extend the PAC learning framework to address learning and control in non-stationary environments. Using this framework, called “Prospective Control”, we prove that under certain fairly general assumptions, empirical risk minimization (ERM) asymptotically achieves the Bayes optimal policy. We then consider a specific instance of prospective control, foraging—which is a canonical task for any mobile agent—be it natural or artificial. We illustrate that existing reinforcement learning algorithms fail to learn in these non-stationary environments, and even with modifications, they are orders of magnitude less efficient than our prospective foraging agents. Code is available at: https://github.com/neurodata/ProspectiveLearningwithControl.
Introduction
When building real-world machine learning and/or artificial intelligence (ML/AI) systems, we almost always care about their future performance because we will be using them in the future. And yet, many classical formulations of ML/AI essentially ignore this fact. Consider, for example, Probably Approximately Correct (PAC) learning [Mohri2012-qj], the primary workhorse of modern ML/AI. The key assumption underlying this framework is that each data point is sampled independently and identically from a given fixed distribution (IID). In other words, the future is assumed to be identical to the past (up to distribution) and also independent of any past data. This assumption is woefully inadequate for many real-world problems.
Reinforcement learning (RL) and online learning (OL) are two popular machine learning paradigms relaxing this IID assumption. Rather than IID, RL typically models the environment as a Markov Decision Process (MDP) with states, actions, rewards, and aims to learn a behavior/policy that maximizes long-term return [Bertsekas2023, Sutton2018-gx, StrehlLiLittman2009PACMDPs]. RL has achieved many real-world successes, e.g. Go [Silver2016NatureGo]. However, standard RL methods fail in non-stationary reward settings and require multiple episodes (though see [Chen2022-ja]). On the other hand, online learning algorithms are run in a sequence of consecutive rounds, making a prediction, observing feedback, and updating [ShalevShwartz2012]. The OL framework tends to make no distributional assumptions about the data whatsoever. This is a worst-case assumption scenario, because it essentially says it knows absolutely nothing about the future, the opposite of IID. Thus, the goal of online learning algorithms is to find the best fixed hypothesis in hindsight.
Neither of these frameworks, therefore, relaxes the IID assumption in a fully satisfactory way. Prospective Learning (PL) [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik] introduced recently generalizes PAC learning to address stochastic processes, rather than independent and identically distributed data. In those prior works, the IID assumption was relaxed, but decisions had no impact on the future. Here, we generalize further to incorporate stochastic decision processes, so that future loss and data distributions also depend on current actions.
As a concrete example, consider foraging. Foraging is informally defined as the sequential search for resources [Barack2024-sx]. It is a central behavior of all mobile agents, be they living organisms, AI bots, or robots. Standard formalizations of foraging have significant limitations in terms of characterizing real-world behaviors. Most importantly for this work, classical models of foraging assume that the environment is essentially stationary. For example, without the single agent’s intervention, resources stay available indefinitely. This is a poor model of real-world foraging because, typically, resources are depleted, rather than persistent.
We therefore introduce prospective learning with control (PL+C), as a generalization to the prior work on prospective learning without control (PL-C) [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik]. We first briefly summarize PL-C using a slightly revised notation (2). Then we formally introduce PL+C (3). Our framework enables us to prove the existence of a sequence of hypotheses whose risks converge to the Bayes optimal risk (4). We then introduce our Prospective Foraging (ProForg) (5). ProForg performs substantially better than classical RL algorithms—even with suitable modifications to address the non-stationarity—on a dynamic foraging problem (7). We therefore believe that PL+C presents a promising strategy for learning to control the future optimally and efficiently.
Prospective Learning without control (PL-C)
Here we briefly review the prior work on this topic, which is called "prospective learning" [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik] (PL), modifying notation slightly for convenience. In retrospect, it would more specifically be called prospective learning without control.
Data Let the input and output be denoted by $x_{t}\in\mathcal{X}$ and $y_{t}\in\mathcal{Y}$, respectively. Let $z_{t}=(x_{t},y_{t})$. We will denote the observed data, $z_{\leq t}=\{z_{1},\dots,z_{t}\}$, and the unobserved data, $z_{>t}$. In contrast to PAC learning, $t$ is not just a dummy variable, but rather, indexes time. Let $t\in\mathcal{T}\subseteq\mathbb{N}$, where $\mathbb{N}=\{1,2,\dots\}$ are the counting numbers, and let $|\mathcal{T}|=T\leq\infty$ be the total number of time steps. We therefore define the data triple $(x_{t},y_{t},t)$.
Hypothesis The hypothesis space generalizes the notion from PAC learning. Here, $h_{t}:\mathcal{X}\times\mathcal{T}\to\mathcal{Y}$ is a hypothesis in $\mathcal{H}_{t}\subseteq\{h:\mathcal{X}\times\mathcal{T}\to\mathcal{Y}\}$, so $h_{t}$ infers an output $y_{t}$ for a given input $x_{t}$ at time $t$.Denote the hypothesis sequence, $h\equiv(h_{1},h_{2},\dots)$, where each element of this sequence $h_{t}:\mathcal{X}\mapsto\mathcal{Y}$. At times, we will consider a sequence of nested hypothesis spaces, $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots\subseteq\mathcal{H}$.
Loss The instantaneous loss, $\bar{\ell}:\mathcal{H}\times\mathbb{Z}\to\mathbb{R}$, is a map from the hypothesis and current data to the reals. In an abuse of notation, we write $\bar{\ell}(h_{t},z_{t})\equiv\bar{\ell}(h(x_{t},t),y_{t})$. In all real-world problems, we care about the integrated future loss. Let $w_{t}$ be a non-increasing non-negative weighting function that sums to one, that is $\sum_{t}w_{t}=1$ and $0\leq w_{t}\leq 1\,\forall t$. We thus define loss at time $t$ as $\ell_{t}(h,z_{>t})=\sum_{s>t}w_{s-t}\bar{\ell}(h(x_{s},s),y_{s})$, which is the weighted cumulative loss over all the future data. [^1] If $T$ is infinite, we replace the sums with limiting sums, e.g., $\lim_{T\to\infty}\sum_{t=1}^{T}w_{t}$, to ensure that the sum converges., [^2] Note that if we let $w_{t}$ be a constant function of $t$, we recover something equivalent to the loss in PAC learning. To fully specify loss, therefore, requires specifying a temporal discounting function, $\bm{w}\equiv(w_{1},w_{2},\dots)$. For example, we could let $\bm{w}$ be an exponentially decaying function, $w_{s-t}=\gamma^{s-t}$, where $\gamma\leq 1$. [^3] This is a slight variation of how prospective loss was previously defined in De-Silva2024-jj.
Learner Let a learner, $\mathcal{L}$ maps from the data to a hypothesis, $\mathcal{L}(\mathcal{D}_{t})=\hat{h}^{t}$, where $\mathcal{D}_{t}=(x_{s},y_{s},s)_{s\leq t}$, and $\hat{h}^{t}$ is the estimated hypothesis sequence obtained by the learner at time tt. [^4] Formally, we assume $h$ is a random variable and $h\in\sigma(Z_{\leq t})$ where $\sigma(\cdot)$ denotes the filtration (an increasing sequence of sigma-algebras) of the stochastic process $Z$.
Assumptions Let $\bm{X}=X_{0<t}$ and $\bm{Y}=Y_{0<t}$ be stochastic processes, taking values $\bm{X}=x_{0<t}$ and $\bm{Y}=y_{0<t}$. Assume $(\bm{X},\bm{Y})\sim F\in\mathcal{F}$. In general, we place no assumptions on $\mathcal{F}$.
Theoretical Goal Define risk as expected loss, $R_{t}(h)=\operatorname*{\mathbb{E}}_{F}[\ell_{t}(h,Z)\mid z_{\leq t}]=\int\ell_{t}(h,Z)d\mathbb{P}_{Z\mid z_{\leq t}}$. [^5] This is a slight abuse of notation, because previously $\ell$ mapped from a hypothesis and data corpus, but now we are saying that it maps from a hypothesis and a collection of random variables. We have used the shorthand $\operatorname*{\mathbb{E}}[Y\mid x]$ for $\operatorname*{\mathbb{E}}[Y\mid X=x]$. Note that $\mathbb{E}[Z|z_{\leq t}]$ is equivalent to $\mathbb{E}[Z_{>t}|z_{\leq t}]$ because the past $z$’s are given. Let $R_{t}^{*}=\min_{h\in\mathcal{H}_{t}}R_{t}(h)$ be the (Bayes) optimal risk. [^6] In a slight abuse of notation, we use $\min$ in place of $\inf$ even if the $\min$ might not exist. Here we use the shorthand $\mathcal{H}_{t}\equiv\sigma(Z_{\leq t})$.
We seek a learner, $\mathcal{L}$ with the following property: for any $F\in\mathcal{F}$, and ${\epsilon},\delta>0$, there exists a $t^{\prime}({\epsilon},\delta)$ such that the learner outputs a hypothesis $\hat{h}^{t}$ satisfying $\operatorname{\mathbb{P}}[R_{t}(\hat{h}^{t})-R_{t}^{*}<{\epsilon}]\geq 1-\delta$, for any $t>t^{\prime}({\epsilon},\delta)$.
One implication of the above desiderata is that the estimated hypothesis, $\hat{h}^{t}$, is a consistent estimator of the risk [Shalev-Shwartz2014-zo]. In other words, $R(\hat{h}^{t})\to R^{*}_{t}$ as $t\to\infty$. The above goal, however, is not an explicit objective function. In practice, therefore, we specify an objective function to minimize. The hope is that if we minimize this objective, then eventually, the resulting estimand satisfies the above desiderata. We seek to minimize the weighted sum of instantaneous losses on future data:
$\hat{h}^{t}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}R_{t}(h)=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}\int\ell_{t}(h,Z)d\mathbb{P}_{Z\mid z_{\leq t}}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}\int\sum_{s>t}w_{s-t}\bar{\ell}(h(x_{s},s),y_{s})d\mathbb{P}_{Z\mid z_{\leq t}}.$ (1)
If $z$ is a discrete set, the above integral becomes a sum. Let $p_{z}$ denote the probability of a particular sequence $z\in\mathbb{Z}$. With this notation, Equation˜1 becomes
$\hat{h}^{t}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}\sum_{z\in\mathbb{Z}:z_{\leq t}}p_{z}\ell_{t}(h,Z)=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}\sum_{z\in\mathbb{Z}:z_{\leq t}}p_{z}\sum_{s>t}w_{s-t}\bar{\ell}(h(x_{s},s),y_{s}).$ (2)
Empirical Goal In practice, in general, we cannot minimize Equation˜1 because it requires solving an intractable integral, and we typically cannot solve Equation˜2, because it requires sampling too many possible futures (there are exponentially many even if $z$ is binary). Moreover, we cannot minimize Equations˜1 and 2 because they both require sampling from an unknown distribution, $\mathbb{P}_{Z\mid z_{\leq t}}$. To address the first issue, we use a Monte Carlo approximate sum, sampling only a subset of possible futures. To address the second issue, rather than sampling from the unobserved future, we sample only from the observed past. With these two approximations in hand, we have
$\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\ell_{t^{\prime}}(h,z_{>t^{\prime}})\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\sum_{s>t^{\prime}}^{t}w_{s-t^{\prime}}\bar{\ell}(h^{t^{\prime}}(x_{s},s),y_{s}).$ (3)
Main result The main result of De-Silva2024-jj is approximating a variant of Equation˜1, under suitable assumptions, satisfies our theoretical goal. The key assumptions are (a) consistency and (b) uniform concentration. The focus in that work was on settings in which our decisions $\hat{h}^{t}(x_{s},s)$ had no impact on future outcomes or distributions. In such cases, we can ignore the sum in Equation˜1, because all that matters for the current decision is the instantaneous loss. Here, we are interested in general settings where the decisions also impact future distributions and losses.
Prospective Learning with Control (PL+C)
We modify the above prospective learning framework to accommodate stochastic decision processes, rather than merely stochastic processes. In other words, we extend the framework from learning to predict the future, to learning to control the future. The key difference is in the loss. For prediction, the loss is the error relative to the truth/optimal, which is always observed. In control, we do not necessarily see the truth/optimal, so one cannot define loss as an error term. Below, we formalize this distinction.
Data Let $x_{t}$ be the position of the agent in an environment at time $t$. Thus, if the environment is two-dimensional, finite, rectangular, and discrete, then it has width $W$ and height $H$, and $x_{t}\in\mathcal{X}\equiv[W]\times[H]$ is the state of the agent at time $t$. As above, we also record time $t\in\mathcal{T}\subset\mathbb{N}$. We refer to the position and time together as the observation state (or simply state), $(x_{t},t)$. Then $y_{t}\in\mathcal{Y}$ is the reward the agent would receive at time $t$ at any of the positions. Thus, if position is a two-dimensional grid as described above, then $\mathcal{Y}\equiv\mathbb{R}_{+}^{W\times H}$. Thus, the past data are the observation state (position and time) and observed rewards.
Hypothesis The hypothesis maps from the current state to the next position, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}$. Without an estimate of the rewards, however, we have no way, in general, to estimate what the rewards would be if we changed the hypothesis. Thus, we could either let the hypothesis map to the next position and the next reward, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}\times\mathbb{R}_{+}$, and so $h(x_{s},s)\to x_{s+1},y_{s+1}(x_{s+1})$. More generally, the hypothesis maps to the next position and the reward at all possible positions, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}\times\mathcal{Y}$, so $h(x_{s},s)\to(x_{s+1},y_{s+1})$.
Loss The instantaneous loss operates on the current hypothesis and data, $\bar{\ell}:\mathcal{H}\times\mathbb{Z}\to\mathbb{R}_{+}$. The instantaneous loss is the negative reward the agent receives at the position the agent is in at that time. Recalling that $x_{s}$ encodes the position of the agent at time $s$, and that $y_{s+1}$ encodes the reward at all locations at time $s+1$, so, in a slight abuse of notation, we can write
$\bar{\ell}(h_{s}(x_{s},s),y_{s+1})=-y_{s+1}(x_{s+1}).$ (4)
The loss at time $t$ is the cumulative instantaneous loss, $\sum_{s>t}w_{s-t}\bar{\ell}(h_{s}(x_{s},s),y_{s+1})$. [^7] We intentionally do not specify whether the horizon is finite or infinite here.
Learner The learner $\mathcal{L}$ is a map from the observed data, $\mathcal{D}_{t}$, to a hypothesis, $\hat{h}^{t}$. The observed data are $\mathcal{D}_{t}=(x_{s},y_{s+1}(x_{s+1}),s)_{s\leq t-1}$.
Assumptions Let $\bm{X}=X_{0,1,2,\dots}$ and $\bm{Y}=Y_{0,1,2,\dots}$ be stochastic decision processes, that is, $(\bm{X},\bm{Y})\sim F\in\mathcal{F}$, where $Y_{>t}$ can be an arbitrary function of $(X_{\leq t},Y_{\leq t})$. Denote the values of the random variables $\bm{X}$ and $\bm{Y}$ as $\bm{x}$ and $\bm{y}$, respectively. Note that the distribution of $X_{t}$ is, in general, a function of the previous sequence of observed $x_{\leq t}$, which is itself a function of the observed rewards $y_{\leq t}$, and the estimated hypotheses, $\hat{h}^{\leq t}$. If we assume that the learners and hypotheses are deterministic functions of the data, then future distributions of $X_{s}$ are functions of the learner and hypotheses, but not conditionally independent on them, once conditioned on the data.
Theoretical Goal The goal of learning a hypothesis whose risk converges to the Bayes optimal risk remains the same in prospective learning with or without control. However, the objective function we seek to optimize is slightly modified because control renders the problem inherently sequential. Thus, the instantaneous loss in PL without control (Equation˜1) is replaced with the instantaneous loss in PL with control (Equation˜4):
$\hat{h}^{t}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}R_{t}(h)\equiv\operatornamewithlimits{argmax}_{h\in\mathcal{H}_{t}}\int\sum_{s>t}w_{s-t}\times y_{s+1}(x_{s+1})d\mathbb{P}_{Z\mid z_{\leq t}}=\operatornamewithlimits{argmax}_{h\in\mathcal{H}_{t}}\sum_{z\in\mathbb{Z}:z_{\leq t}}p_{z}\sum_{s>t}w_{s-t}\times y_{s+1}(x_{s+1}).$ (5)
Note that while the $\operatornamewithlimits{argmax}$ is with respect to $h$, on the right-hand side after the $\equiv"$, the $h$ is implicit. We could notationally let $y_{s}$ be a function of $h$ to clarify if we desired.
Empirical Goal Much like PL without control, in PL with control, we cannot directly solve Equation˜5 because we cannot sample all possible futures, and we do not know the likelihood of each possible future. Thus, we modify the sums in Equation˜5 to only sum up to the current time, $t$:
$\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\sum_{s>t^{\prime}}^{t}-w_{s-t^{\prime}}\times y_{s+1}(x_{s+1}).$ (6)
Relationship to prospective learning without control The observed data triples in PL+C are $(x_{s},y_{s+1}(x_{s+1}),s)$, rather than $(x_{s},y_{s},s)$ in PL-C. Note that these data triples differ in two ways. First, the index on $y$ changes from $s$ to $s+1$ as we go from PL-C to PL+C; this is required because the rewards are necessarily a function of the action of the agent. Second, rather than the entire vector of output being observed as in PL-C, in PL+C, we only obtain a partial observation: the reward at the location we happen to currently inhabit. This renders PL+C closely related to causal inference, in which the outcomes are defined as a function of the intervention (or treatment), and therefore only partially observed, leading to the potential outcomes framework [Neyman1923-aw, Rubin1974-gt]. Hypotheses are no longer maps from $\mathcal{X}\times\mathcal{T}$ to $\mathcal{Y}$, rather, they map to $\mathcal{X}$ and $\mathbb{R}_{+}$. Loss in PL with or without control is the cumulative instantaneous loss. However, in PL-C, instantaneous loss is typically an error signal, e.g., the difference between the predicted output and the actual output. In PL+C, loss is just a negative reward. Moreover, loss is not observed for any of the decisions that the agent did not make, necessitating the use of some surrogate loss function in practice. Learners map from data to hypotheses, though the data are subtly different due to the time indices and partial observation. Assumptions can be the same, though note that the future data are now necessarily a function of the learner and hypothesis in PL+C, whereas they were not necessarily in PL-C. The theoretical goal is the same, modulo replacing the loss from PL-C with that from PL+C. The empirical goal changes somewhat similarly, insofar as in PL+C we cannot directly solve Equation˜5, so we use a surrogate optimization problem, Equation˜6.
Theoretical Results
Here we provide proof sketches; formal proofs are in Appendix˜A.
Lemma 4.1.
Suppose $\{Z_{t}\}_{t=1}^{\infty}$ is a stochastic process and let $\{\mathcal{H}_{t}\}_{t=1}^{\infty}$ be an increasing hypothesis class, such that there exists $h^{(t)}\in\mathcal{H}_{t}$ which satisfies $\lim_{t\to\infty}\mathbb{E}[R_{t}(h^{(t)})-R_{t}^{*}]=0.$ Additionally, suppose $\{u_{t}\}_{t=1}^{\infty}$ is a sequence such that $u_{t}\to\infty$ as $t\to\infty$, and $u_{t}\leq t$ for all $t$. Define the partial cumulative prospective loss to be $e_{m}(h)=\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1}).$
Then, there exists $\tilde{h}^{(t)}\in\mathcal{H}_{t}$ such that $\lim_{t\to\infty}\mathbb{E}\left[\sup_{u_{t}\leq m\leq\infty}e_{m}(\tilde{h}^{(t)})|Z_{\leq t}\right]-R^{*}_{t}=0.$
Sketch of proof. A condition for the above lemma to hold is the existence of a sequence of hypotheses whose risks converge to the Bayes optimal risk. We establish that for this asymptotically optimal sequence of hypotheses, there exists a subsequence of timepoints on which the expected partial cumulative loss is arbitrarily close to the Bayes risk, using Markov’s Inequality and Borel-Cantelli Lemma. We use this subsequence of timepoints to establish that for a particular sequence of hypotheses, the expected partial cumulative loss approaches the Bayes optimal risk.
Theorem 4.1.
Consider a family of stochastic processes $(\mathbf{X}_{t},\mathbf{Y}_{t})_{t>0}$. Suppose there exists a hypothesis class such that $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\dots$ such that $\lim_{t\to\infty}\left[\inf_{h\in\mathcal{H}_{t}}R_{t}(h)-R^{*}_{t}\right]=0.$ Also, there exist sequences $\{u_{t}\}_{t=1}^{\infty}$ and $\{\xi_{t}\}_{t=1}^{\infty}$ satisfying $\lim_{t\to\infty}\xi_{t}=0$, and $u_{t}\leq t$, such that
$\mathbb{E}\left[\max_{h\in\mathcal{H}_{t}}\left|\sum_{s=t+1}^{\infty}w_{s-t}\tilde{y}_{s+1}(x_{s+1})-\max_{u_{t}\leq m\leq t}\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1})\right|\right]\leq\xi_{t}.$
Then, there exists a sequence $\{i_{t}\}_{t=1}^{\infty}$ such that
$\hat{h}^{(t)}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{i_{t}}}\max_{u_{i_{t}}\leq m\leq t}\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1})$~ reaches the Bayes optimal risk.
Sketch of proof. We note that for any sequence of hypotheses, the difference between the prospective loss and the partial cumulative loss is bounded by terms approaching zero. We find a subsequence of timepoints on which this difference is sufficiently small, and using Lemma 4.1 we establish that the Bayes risk of the minimizer of the partial cumulative loss on that particular subsequence of timepoints, approaches zero.
Estimation & inference for prospective learning with control
Inspired by online re-planning in RL [Bertsekas2023], we adopt an online strategy here. Before iterative policy updates, we start with a short warm-up phase. That is, the agent follows a nearly random policy to gather initial interaction data (states, rewards), after which we fit an initial model. This pretraining alleviates the cold start and enables more effective exploration at the beginning of the sequence. After the pretraining phase, at each online iteration, we estimate the cumulative loss and the instantaneous loss, and then decide the next step based on the currently updated learners.
Warm Start via Batch Pretraining
Cumulative loss estimation
Recall the empirical goal in prospective learning with control, Equation˜6. Based on this equation, we build a data corpus using prospective loss as follows. Let $\widetilde{x}_{s}\in\mathbb{R}^{d_{p}}$ and $\widetilde{s}\in\mathbb{R}^{d_{t}}$ be the embedded positions and times, and let $d=d_{p}+d_{t}$. Let $X_{t}=(\widetilde{x}_{s},\widetilde{s})_{s\leq t}$ be the input matrix, such that $X_{t}\in\mathbb{R}^{d\times t}$. Let $Y_{t}\in\mathbb{R}^{t}$ be the sequence of cumulative rewards, that is $Y_{t}=(\sum_{t^{\prime}=s+1}^{t}-w_{s-t^{\prime}}\times y_{t}(x_{t^{\prime}}))_{s\leq t}$. Then, we can feed data into any supervised learning machine algorithm, for example, a decision forest or a deep network. Let the resulting regression function be denoted $\hat{g}^{p}$, which maps from a state, $(x_{s},s)$ to an estimated prospective loss $\bar{y}_{s+1}\in\mathbb{R}_{+}$.  [^8] We implicitly assume that $\hat{g}$ implements the embedding function.
Instantaneous loss estimation
When searching over hypotheses, some will result in positions at various times that are not identical (i.e., counterfactual) to the positions that were selected by the estimated hypothesis at that time, $\hat{h}^{s}$. We therefore lack observations of $y_{s}$ for all the counterfactual positions, and in practice, we cannot actually solve Equation˜6, because we have missing data. Instead, we introduce a further approximation, replacing $y_{s+1}$ with $\tilde{y}_{s+1}$. Define $\widetilde{y}_{s}$ as follows. Let $x^{\prime}_{s}$ denote potential outcomes that the agent can take at time $s$, whereas $x_{s}$ is the position that the agent actually took at that time, and let $\hat{y}_{s}(x^{\prime}_{s})$ denote the estimated reward at time $s$ at position $x^{\prime}_{s}$ based on the current hypothesis:
$\widetilde{y}_{s}(x_{s})=\begin{cases}y_{s}(x_{s})&x^{\prime}_{s}=x_{s}\\ \hat{y}_{s}(x^{\prime}_{s})&x^{\prime}_{s}\neq x_{s}.\qquad\triangleright\text{these are counterfactuals}\end{cases}$ (7)
Putting all this together yields: $\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\sum_{s>t^{\prime}}^{t}-w_{s-t^{\prime}}\times\tilde{y}_{s+1}(x_{s+1}).$
Since we are estimating $y$ at this point, we are no longer restricted to summing up until time $t$ for the inner sum. Let $T$ be some predefined horizon, and we have $\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\sum_{s>t^{\prime}}^{t+T}-w_{s-t^{\prime}}\times\tilde{y}_{s+1}(x_{s+1}).$ In this case, we can build a data corpus as follows. $X_{t}$ remains the same as it is above. However, $Y_{t}=(\widetilde{y}_{s+1}(x^{\prime}_{s+1}))_{s\leq t}$, which is also in $\mathbb{R}^{t}$. Let the resulting regression function be denoted $\hat{g}^{i}$, which maps from a state $({x}_{s},{s})$ to an estimated instantaneous loss $\hat{y}\in\mathbb{R}_{+}$.
Online estimation and inference
Recall that our goal is Equation˜6. And now, we get estimated instantaneous reward $\tilde{y}_{s}$, and cumulative reward $\bar{y}_{s}$, respectively. We combine these two parts to approximate the infinite horizon prospective loss. This yields both counterfactual insight and a tractable approximation to future loss. To be more specific, given a predefined horizon, we have $\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\{w_{t+T-t^{\prime}}\times\bar{y}_{t+T}+\sum_{s>t^{\prime}}^{t+T}-w_{s-t^{\prime}}\times\tilde{y}_{s+1}(x_{s+1})\}$, where $\bar{y}_{t+T}$ represents an estimation of terminal cost [Bertsekas2023] at time $t+T$, and $\tilde{y}_{s}$ is the estimated instantaneous cost at time $s$. Here, two supervised learning models are used: one for approximating the look-ahead cost for the finite steps, and another to approximate the terminal cost correcting the truncation of large or infinite horizons.
Inference in Practice
We evaluate each potential future trajectory via calculating its prospective loss, $\{w_{T}\times\hat{g}^{p}(x_{s+T},s+T)+\sum_{j=s}^{s+T}w_{j-s}\times\hat{g}^{i}(x_{s},s)\}$. Assuming $T=6$, and $A=2$, we have $2^{6}=64$ different trajectories, each with an associated prospective loss, as estimated by iterating the instantaneous losses and adding the terminal cost at the end. We select the next state that has the trajectory with the maximum value. For each time $t$, the decision is made by the learner trained on past observations till time t1t-1. Algorithm 1 in Appendix 6 provides details. The above is an exhaustive search strategy. If our computational budget is inadequate to search all the possible trajectories, then one can search a subset of possible trajectories using Monte Carlo Tree Search (MCTS) [Bruegmann1993mcgo, Kocsis2006uct, Coulom2007mcts, Silver2016NatureGo, silver2017masteringchessshogiselfplay].
Algorithm
Algorithm 1 ProForg: Online learning to control the future in non-stationary environments.
1:Observed data triples $\{(x_{s},s,y_{s+1}(x_{s}))\}_{s=1}^{T}$, discount $\gamma\in[0,1)$, horizon $H$, embedding function $\varphi(\cdot)$ , learning algorithm $\mathcal{A}(\cdot)$,
[…]
3:for $s=1$ to $T$
4:  $x^{\prime}_{s}\leftarrow\varphi(x_{s},s)$ $\triangleright$ embed position and time
5:  $y^{\prime}_{s}\leftarrow\sum_{t^{\prime}=s+1}^{T}\gamma^{\,t^{\prime}-(s+1)}\,y_{t^{\prime}}$ $\triangleright$ discounted future return starting at s+1s{+}1
6:end for
[…]
9:Online updates $\hat{g}^{i}$ and $\hat{g}^{p}$ and $x_{s}$
10:for $s=1>T$
11:  $\mathcal{C}\leftarrow\textsc{AllPaths}(x_{s},H)$ $\triangleright$ each $x^{(s:H+s)}\in\mathcal{C}$ is a length-$H$ trajectory, $(x^{(s)},\dots,x^{(H+s)})$
12:  for $x^{(s:H+s)}\in\mathcal{C}$
13:  $\hat{G}(x^{(s:H+s})\leftarrow\sum_{h=1}^{H}\gamma^{\,h-1}\;\mathrm{\hat{g}^{i}}(\varphi(x^{(h)},s{+}h))$ $\triangleright$ approx. finite horizon loss
14:  $\hat{G}(x^{(s:H+s)})\leftarrow\hat{G}(x^{(s:H+s)})+\gamma^{H}\hat{g}^{p}(\varphi(x^{(H)},s{+}H))$ $\triangleright$ approx. terminal loss
15:  end for
16:  $x^{(1:H)}_{\star}\leftarrow\arg\max_{x^{(1:H)}\in\mathcal{C}}\hat{G}(x^{(1:H)})$
17:  $x_{s+1}\leftarrow x_{\star}^{(1)}$ $\triangleright$ execute only the first step of the best path
18:  $x^{\prime}_{s+1}\leftarrow\varphi(x_{s+1},s+1)$
19:  $\hat{g}^{i}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{i}_{s},x^{\prime}_{s+1},y_{s+1}\right)$ $\triangleright$ update regressor on instantaneous loss
20:  $\hat{g}^{p}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{p}_{s},x^{\prime}_{s+1},y^{\prime}_{s+1}\right)$ $\triangleright$ update regressor on cumulative loss
21:end for
22:return $\hat{g}^{i},\hat{g}^{p},x_{s+1}$ $\triangleright$ updated regressors and states
23:
Experimental Validation
To validate the above-described theoretical framework, we devise a specific prospective learning with control problem, namely, prospective foraging.
Prospective foraging
Informally, we assume an agent is navigating an environment, and resources (also called affordances [Gibson2014-or] or rewards [Sutton2018-gx]) appear and disappear in that environment at various locations. At any given time, the agent may move its position, and the loss is the negative reward associated with the agent’s current location.
Data Assume that $\mathcal{X}$ is a discrete two-dimensional grid, and let $H=1$, so the environment is a chain (Figure 1). Thus, $x_{s}\in\{1,\dots,W\}$, and $y_{s}\in\mathbb{R}_{+}^{W}$. We potentially encode position via a one-hot encoding, thereby $x_{s}$ becomes a binary $W$-dimensional vector. We also embed time into a $50$-dimensional vector using the classical ‘positional’ encoding strategy, as in De-Silva2024-jj. Let $\tilde{x}_{s}$ and $\tilde{s}$ represent the embedded position and time.
Hypothesis At each time step, the hypothesis can move the agent by one cell in the grid in either direction, assuming it is not at a boundary, or stay put. At a boundary, it can only move nowhere or in the opposite direction. Formally, $h(x_{s},s)\to x_{s+1}\in\mathcal{X}_{s}$, where $\mathcal{X}_{s}$ could be either $\{x_{s}-1,x_{s},x_{s}+1\}$, $\{x_{s}-1,x_{s}\}$, or $\{x_{s},x_{s}+1\}$.
Loss The loss is defined as above, with the temporal discounting factor of $\gamma$.
Assumptions Let $F_{\bm{Y}}$ denote the prior distribution of $\bm{Y}$. Assume that $\bm{Y}$ is independent of $\bm{X}$, i.e. $F_{\bm{Y}|\bm{X}}=F_{\bm{Y}}$. Let $y_{s}\in\mathbb{R}_{+}^{W}$ be a $W$-dimensional non-negative vector, where each element encodes the reward and the agent would receive at location $x_{s}$ at time $s$. Assume that for two locations, $A$ and $B$, there are rewards sometimes, but for the other W2W-2 locations, there are never any rewards. For locations $A$ and $B$, assume that rewards arrive on a schedule. At every $r$ time steps, location $A$ receives a boost of one reward, which then decays exponentially down to zero at a rate of $\tau$. Location $B$ has the same temporal dynamics, but shifted by $r/2$ time steps. Thus, $y_{s}(A)=e^{-(s\mod r)/\tau}$, $y_{s}(B)=e^{-((s+r/2)\mod r)/\tau}$, where $s\mod r$ indicates the modulus function. The dynamics of the conditional distribution, $F_{\bm{X}|\bm{Y}}$, are a function of the learned hypotheses, $\hat{h}$.
Figure 1: 1-D foraging environment. An agent moves along a $1\times 7$ linear track with two reward patches (A, B). Rewards alternate between the two patches over time, and the currently active patch’s reward decays exponentially.
We implement and evaluate the prospective forager (ProForg) within the 1-D foraging environment described in Section˜7.1 and Bai2026-ik: the agent forages along a $1\times 7$ linear track containing only two reward patches, A and B, located three grid cells apart, as shown in Figure˜1. Every $r=10$ timestamps, rewards alternatively appear at A and B. The agent’s goal is to maximize the total reward in its single lifetime, meaning there are no resets within each run.
We define the prospective regret of a hypothesis $h^{(t)}$ as the difference between the prospective risk returned by hypothesis $h^{(t)}$ and Bayes optimal risk: $\Delta{h^{(t)}}=R_{t}(h^{(t)})-R_{t}^{*}$. In practice, one cannot evaluate regret over an infinite future. Therefore, we approximate the prospective regret by evaluating it over a finite horizon of $T$ time steps and returning the normalized one (NPR): $\frac{1}{T}\Big(R_{t,T}\!\big(h^{(t)}\big)-R_{t,T}^{\star}\Big)$, where $R_{t,T}$ is the finite horizon prospective risk. In addition, since we only have one trajectory after time t, we replace the risk by the prospective loss, so $\Delta{h^{(t)}}\approx\frac{1}{T}\sum_{i=s}^{s+T}w_{i-s}(y_{s}(x^{*}_{s})-y_{s}(x_{s}))$, where $(x^{*}_{s},x^{*}_{s+1},....,x^{*}_{s+T})$ are the states returned by Bayes optimal hypothesis at the corresponding times.
Numerical Results
ProForg
After an initial warm-up phase, we let the ProForg learn in an online manner as it interacts with the environment over time (Figure˜2). The procedure follows the steps detailed in Algorithm˜1. During the online updating part, we retrained regressors with all data collected so far for each update. We compute the normalized prospective regret of ProForg over time, and compare it to Fitted Q-Iteration (FQI) [ernst2005tree, munos2008finite] and Soft Actor-Critic (SAC) [haarnoja2018soft, haarnoja2018softa], two classical reinforcement learning algorithms, and their time-aware variants that we introduce here to enable them to work in non-stationary environments (see Appendices˜B, 2 and 3). In both ProForg and FQI variants, we use Random Forests as our chosen class of regressors.
Figure 2: ProForg efficiently achieves Bayes optimal regret. Normalized prospective regret of ProForg (red), time-aware Fitted Q-Iteration (FQI with time, blue-purple, our invention to improve FQI), Time-agnostic Fitted Q-Iteration(FQI w/o time, light-blue [ernst2005tree]), time-aware Soft Actor-Critic (SAC with time, purple-red), and Time-agnostic Soft Actor-Critic(SAC w/o time, lavender [haarnoja2018soft]). While ProForg, time-aware FQI, and time-aware SAC converge to having zero regret, ProForg is orders of magnitude more efficient than either of them. And time-agnostic variants converge to a sub-optimal regret regardless of the time spent interaction.
Online or Offline?
Building on the online formulation, we compare the online and offline ProForg, under the same environment settings and parameters for training and inference (Figure˜3). The only difference is the next state selection: online-ProForg chooses the next step using its currently learned policy, while offline-ProForg explores the next step uniformly at random.
[…]
Instantaneous, Cumulative or Both?
ProForg uses two learners during training and inference phases. $\mathrm{ML_{I}}$, fit for instantaneous loss, and $\mathrm{ML_{C}}$ for future cumulative loss. During inference, we combine a finite-horizon “lookahead” from $\mathrm{ML_{I}}$ with a terminal cost approximation from $\mathrm{ML_{C}}$. We compare offline invariants of Instantaneous-only (ProForg-I), Future-only (ProForg-C), and combined (ProForg) (Figure˜4). Using the terminal cost alone (ProForg-C) converges inefficiently, but we find adding it to the instantaneous loss accelerates learning versus instantaneous-only.
Figure 4: Normalized prospective regret for ProForg(red), ProForg-I (orange), and ProForg-C (yellow).Removing either component reduces performance relative to ProForg.
ProForg with Neural Network
In Figure˜5, we show that neural networks can also be used as the regressor in ProForg. While the ProForg with a neural network (specifically, a multi-layer perceptron with two hidden layers comprising 128 units) converges eventually, the convergence is about an order of magnitude slower than the ProForg employing a random forest as the regressor.
Figure 5: ProForg with decision forests is 4x more efficient than with neural networks. Normalized prospective regret for ProForg with Gradient-Boosted Trees (red) and MLP Regressor (blue). While ProForg is 4x more efficient, ProForg-NN does converge as well.
Discussion
In this work, we introduce the framework of prospective learning with control. Prospective Foraging, a concrete example of prospective learning with control, learns to optimally behave in the future in a dynamic environment. Our results show that our algorithm, ProForg, can learn to plan ahead optimally. We also observe that standard RL baselines fail to converge to the optimal policy despite the simple nature of the one-dimensional foraging environment. Though much work remains to be done, we believe prospective learning with control is an initial step towards understanding and utilizing optimal control of the future in both natural and artificial intelligences. We leave generalizing to continuous action, state, and time spaces, and deploying our algorithms in complex real-world environments, as future work.
Appendix A Theoretical Results
Lemma A.1.
Suppose $\{Z_{t}\}_{t=1}^{\infty}$ is a stochastic process and let $\{\mathcal{H}_{t}\}_{t=1}^{\infty}$ be an increasing hypothesis class, such that there exists $h^{(t)}\in\mathcal{H}_{t}$ which satisfies
$\lim_{t\to\infty}\mathbb{E}[R_{t}(h^{(t)})-R_{t}^{*}]=0.$ (8)
Additionally, suppose $\{u_{t}\}_{t=1}^{\infty}$ is a sequence such that $u_{t}\to\infty$ as $t\to\infty$, and $u_{t}\leq t$ for all $t$. Define
$e_{m}(h)=\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1}).$
Then, there exists $\tilde{h}^{(t)}\in\mathcal{H}_{t}$ such that
$\lim_{t\to\infty}\mathbb{E}\left[\sup_{u_{t}\leq m\leq\infty}e_{m}(\tilde{h}^{(t)})|Z_{\leq t}\right]-R^{*}_{t}=0.$
Proof. There exists $h^{(t)}\in\mathcal{H}_{t}$ such that
$\lim_{t\to\infty}\mathbb{E}[R_{t}(h^{(t)})-R_{t}^{*}]=0.$
Note that $R_{t}(h)\geq R_{t}^{*}$ almost surely. Choose a subsequence $\{j_{k}\}_{k=1}^{\infty}$ such that
$\mathbb{E}[R_{j_{k}}(h^{(j_{k})})-R_{t}^{*}]\leq\frac{1}{4^{k}}.$
Now,
$\mathbb{E}[R_{t}(h^{(t)})-R_{t}^{*}]$ $=\mathbb{E}\left[\mathbb{E}[\lim\sup_{m\to\infty}e_{m}(h^{(t)})|Z_{\leq t}]-R_{t}^{*}\right]$
$=\mathbb{E}\left[\mathbb{E}[\lim_{i\to\infty}\sup_{u_{i}\leq m\leq\infty}e_{m}(h^{(t)})|Z_{\leq t}]-R_{t}^{*}\right]$
$=\lim_{i\to\infty}\mathbb{E}\left[\mathbb{E}[\sup_{u_{i}\leq m\leq\infty}e_{m}(h^{(t)})|Z_{\leq t}]-R_{t}^{*}\right]$
Thus, for every $k$, there exists integer $i_{k}$ such that
$\mathbb{E}\left[\mathbb{E}[\sup_{u_{i_{k}}\leq m\leq\infty}e_{m}(h^{(j_{k})})|Z_{\leq j_{k}}]-R_{j_{k}}^{*}\right]\leq\mathbb{E}[R_{j_{k}}(h^{(j_{k})})-R_{j_{k}}^{*}]+\frac{1}{4^{k}}\leq\frac{2}{4^{k}}.$
Note that,
$\mathbb{E}\left[\sup_{u_{i}\leq m\leq\infty}e_{m}(h^{(t)})|Z_{\leq t}\right]\geq\mathbb{E}\left[\sum_{s=t+1}^{m}w_{s-t}\tilde{y}_{s+1}(x_{s+1})|Z_{\leq t}\right]\geq R_{t}^{*}.$
Using Markov’s Inequality,
$\sum_{k=0}^{\infty}\mathbb{P}\left(\mathbb{E}\left[\sup_{u_{i_{k}}\leq m\leq\infty}e_{m}(h^{(j_{k})})|Z_{\leq j_{k}}\right]-R_{j_{k}}^{*}>2^{\frac{1}{2}-k}\right)$
$\leq\sum_{k=0}^{\infty}\frac{1}{2^{\frac{1}{2}-k}}\mathbb{E}\left[\mathbb{E}\left[\sup_{u_{i_{k}}\leq m\leq\infty}e_{m}(h^{(j_{k})})|Z_{\leq j_{k}}\right]-R_{j_{k}}^{*}\right]$
$\leq\sum_{k=0}^{\infty}\frac{1}{2^{\frac{1}{2}-k}}\frac{2}{4^{k}}$
$<\infty.$
Thus, by Borel-Cantelli Lemma, there exists an integer $k_{0}$ such that for all $k\geq k_{0}$,
$\mathbb{E}\left[\sup_{u_{i_{k}}\leq m\leq\infty}e_{m}(h^{(j_{k})})|Z_{\leq j_{k}}\right]-R_{j_{k}}^{*}\leq 2^{\frac{1}{2}-k}.$
Now, we define $k_{t}=\max\{k\in\mathbb{N}\cup\{0\}:\max\{i_{k},j_{k}\}\leq t\}$. Note that $k_{t}\to\infty$ as $t\to\infty$, because $i_{k},j_{k}\leq\infty$. Define $\alpha_{t}=2^{\frac{1}{2}-k_{t}}$, and observe that $\lim_{t\to\infty}\alpha_{t}=0$. Let $t_{0}\in\mathbb{N}$ be a random variable such that for all $t\geq t_{0}$, $k_{t}\geq k_{0}$, and thus,
$\mathbb{E}\left[\sup_{u_{i_{k_{t}}}\leq m\leq\infty}e_{m}(h^{(j_{k_{t}})})\big|Z_{\leq j_{k_{t}}}\right]-R^{*}_{j_{k_{t}}}\leq\alpha_{t}.$
Choose $\tilde{h}^{(t)}=h^{(j_{k_{t}})}$ for all $t$. Note that $j_{k_{t}}\leq t\Rightarrow\mathcal{H}_{j_{k_{t}}}\subset\mathcal{H}_{t}$. Hence, for $t\geq t_{0}$,
$\mathbb{E}\left[\sup_{u_{t}\leq m\leq\infty}e_{m}(\tilde{h}^{(t)})|Z_{\leq j_{k_{t}}}\right]-R_{j_{k_{t}}}^{*}\leq\mathbb{E}\left[\sup_{u_{i_{k_{t}}}\leq m\leq\infty}e_{m}(\tilde{h}^{(t)})|Z_{\leq j_{k_{t}}}\right]-R^{*}_{j_{k_{t}}}\leq\alpha_{t}.$
Since $\alpha_{t}\to 0$ as $t\to\infty$, we have,
$\lim_{t\to\infty}\bigg(\mathbb{E}\left[\sup_{u_{t}\leq m\leq\infty}e_{m}(\tilde{h}^{(t)})|Z_{\leq j_{k_{t}}}\right]-R^{*}_{j_{k_{t}}}\bigg)=0.$
Theorem A.1.
Consider a family of stochastic processes $(\mathbf{X}_{t},\mathbf{Y}_{t})_{t>0}$. Suppose there exists a hypothesis class such that $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\dots$ such that
$\lim_{t\to\infty}\left[\inf_{h\in\mathcal{H}_{t}}R_{t}(h)-R^{*}_{t}\right]=0.$
Also, there exist sequences $\{u_{t}\}_{t=1}^{\infty}$ and $\{\xi_{t}\}_{t=1}^{\infty}$ satisfying $\lim_{t\to\infty}\xi_{t}=0$, and $u_{t}\leq t$, such that
$\mathbb{E}\left[\max_{h\in\mathcal{H}_{t}}\left|\sum_{s=t+1}^{\infty}w_{s-t}\tilde{y}_{s+1}(x_{s+1})-\max_{u_{t}\leq m\leq t}\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1})\right|\right]\leq\xi_{t}.$
Then, there exists a sequence $\{i_{t}\}_{t=1}^{\infty}$ such that
$\hat{h}^{(t)}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{i_{t}}}\max_{u_{i_{t}}\leq m\leq t}\sum_{s=1}^{m}w_{s}\tilde{y}_{s+1}(x_{s+1})$
reaches the Bayes’ optimal risk.
Proof. We consider a subsequence $\{i_{t}\}_{t=1}^{\infty}$ such that $\xi_{i_{t}}\to 0$ exponentially so that $\sum_{t=1}^{\infty}\xi_{i_{t}}<\infty$. From Markov’s Inequality,
$\sum_{t=0}^{\infty}\mathbb{P}\left[\max_{h\in\mathcal{H}_{i_{t}}}\left|\bar{l}_{t}(h,Z)-\max_{u_{i_{t}}\leq m\leq i_{t}}e_{m}(h)\right|>\sqrt{\xi_{i_{t}}}\right]$
$\leq\sum_{t=0}^{\infty}\frac{1}{\sqrt{\xi_{i_{t}}}}\mathbb{E}\left[\max_{h\in\mathcal{H}_{i_{t}}}\left|\bar{l}_{t}(h,Z)-\max_{u_{i_{t}}\leq m\leq i_{t}}e_{m}(h)\right|\right]$
$\leq\sum_{t=0}^{\infty}\sqrt{\xi_{i_{t}}}$
$<\infty.$
By the Borel-Cantelli Lemma, there exists $t_{1}\in\mathbb{N}$, such that for all $t\geq t_{1}$,
$\max_{h\in\mathcal{H}_{i_{t}}}\left|\bar{l}_{t}(h,Z)-\max_{u_{i_{t}}\leq m\leq i_{t}}e_{m}(h)\right|\leq\sqrt{\xi_{i_{t}}}.$
Let $j_{t}=\max\{t^{\prime}:i_{t^{\prime}}\leq t\}$. Observe that $i_{j_{t}}\to\infty$ as $t\to\infty$. Since $i_{j_{t}}\leq t$,
$\bar{l}_{t}(h,Z)-\max_{u_{i_{j_{t}}}\leq m\leq t}e_{m}(h)\leq\bar{l}_{t}(h,Z)-\max_{u_{i_{j_{t}}}\leq m\leq i_{j_{t}}}e_{m}(h).$
Construct a random variable $t_{2}$ such that for all $t\geq t_{2}$$j_{t}\geq t_{1}$. Thus, for all $t\geq t_{2}$,
$\max_{h\in\mathcal{H}_{i_{j_{t}}}}\left\{\bar{l}_{t}(h,Z)-\max_{u_{i_{j_{t}}}\leq m\leq t}e_{m}(h)\right\}$
$\leq\max_{h\in\mathcal{H}_{i_{j_{t}}}}\left|\bar{l}_{t}(h,Z)-\max_{u_{i_{j_{t}}}\leq m\leq i_{j_{t}}}e_{m}(h)\right|$
$\leq\sqrt{\xi_{i_{j_{t}}}}.$
Choose $i_{t}=i_{j_{t}}$ and notice that since $i_{t}\to\infty$, we have $i_{t}\leq t$. With probability one, for all $t\geq t_{2}$,
$\max_{h\in\mathcal{H}_{i_{t}}}\left\{\bar{l}_{t}(h,Z)-\max_{u_{i_{t}}\leq m\leq t}e_{m}(h)\right\}\leq\sqrt{\xi_{i_{t}}}.$
For $t\geq t_{2}$,
$\bar{l}_{t}(\hat{h}^{(t)},Z)$ $\leq\max_{u_{i_{t}}\leq m\leq t}e_{m}(\hat{h}^{(t)})+\sqrt{\xi_{i_{t}}}$
$\leq\max_{u_{i_{t}}\leq m\leq t}e_{m}(h^{(t)})+\sqrt{\xi_{i_{t}}}$
$\leq\sup_{u_{i_{t}}\leq m\leq t}e_{m}(h^{(t)})+\sqrt{\xi_{i_{t}}}$
Hence,
$\mathbb{E}[\bar{l}_{t}(\hat{h}^{(t)},Z)|Z_{\leq t}]-\mathbb{E}\left[\sup_{u_{i_{t}}\leq m\leq t}e_{m}(h^{(t)})|Z_{\leq t}\right]\to 0$
almost surely. By Lemma 3, we have,
$\mathbb{E}\left[\bar{l}_{t}(\hat{h}^{(t)},Z)|Z_{\leq t}\right]-R_{t}^{*}\to 0.$
Hence, by the bounded convergence theorem,
$0=\mathbb{E}\left[\lim_{t\to\infty}\left(\mathbb{E}\left[\bar{l}_{t}(\hat{h}^{(t)},Z)|Z_{\leq t}\right]-R_{t}^{*}\right)\right]=\lim_{t\to\infty}\mathbb{E}\left[\mathbb{E}\left[\bar{l}(\hat{h}^{(t)},Z)|Z_{\leq t}\right]-R_{t}^{*}\right].$
Appendix B Reinforcement Learning Baselines
Let $s_{t}\in\mathcal{S},a_{t}\in\mathcal{A}$ and $r_{t}\in\mathbb{R}$ denote the state, action, and reward at time $t$, respectively.
Fitted Q-Iteration
Suppose we are given a dataset of interactions $\mathcal{D}=\{(s_{t},a_{t},r_{t})\}_{t=1}^{T}$ collected using some stochastic behavior policy by interacting with a Markov Decision Process (MDP). The generic Fitted Q-iteration (FQI) [ernst2005tree, munos2008finite] algorithm approximates the state-action value function $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ by running the following iterative condition. At iteration $k$, it constructs training targets
$y_{t}^{(k)}=r_{t}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q_{k}(s_{t+1},a^{\prime}),$
from the dataset $\mathcal{D}$, and then it updates $Q_{k+1}$ by fitting a regressor from a model class $\mathcal{F}$ onto these targets.
$Q_{k+1}=\operatornamewithlimits{argmin}_{f\in\mathcal{F}}\sum_{t=1}^{T}\left(f(s_{t},a_{t})-y_{t}^{(k)}\right)^{2}$
Once we have the approximated $\hat{Q}-$function, we use the greedy policy given by,
$\pi(s)=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\hat{Q}(s,a)$
In the time-aware variant of FQI, we make the $Q$-function a function of time $t$ in addition to the state $s$ and action $a$. We do this by slightly modifying the iterative condition as follows. At iteration $k$, it constructs training targets
$y_{t}^{(k)}=r_{t}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q_{k}(s_{t+1},a^{\prime},t),$
from the dataset $\mathcal{D}$, and then it updates $Q_{k+1}$ by fitting a regressor from a model class $\mathcal{F}$ onto these targets.
$Q_{k+1}=\operatornamewithlimits{argmin}_{f\in\mathcal{F}}\sum_{t=1}^{T}\left(f(s_{t},a_{t},t)-y_{t}^{(k)}\right)^{2}$
As before, once we have the approximated $\hat{Q}-$function, we construct a greedy policy by,
$\pi(s,t)=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\hat{Q}(s,a,t)$
However, unlike the static policy from generic FQI, the time-aware variant leads to a dynamic policy that varies with time. In our experiments, we use a Random Forest with $1000$ trees as the regressor of the FQI. We illustrate the pseudocode of online FQI in algorithm˜2.
Algorithm 2 Online Fitted Q-Iteration with $\epsilon$-greedy exploration
1:Initialize Foraging environment env, experience buffer $\mathcal{D}$, warmup size $W$, update interval II, exploration parameter $\epsilon$, action space $\mathcal{A}=\{-1,0,1\}$
[…]
3:for $t=1$ to $T$
4:  $s_{t}\leftarrow\texttt{env}.\text{get\_current\_state()}$
5:  Sample $u\sim\text{Uniform}[0,1]$
6:  if $u<\max(0.01,\epsilon\cdot 0.999^{t})$ or $Q$ is None
7:    $a_{t}\leftarrow\text{RandomChoice}(\mathcal{A})$ $\triangleright$ Explore
8:  else
9:    $a_{t}\leftarrow\operatornamewithlimits{argmax}_{a^{\prime}\in\mathcal{A}}\hat{Q}(s,a^{\prime},t)$ $\triangleright$ Exploit
10:  end if
11:  $(s_{t+1},r_{t})\leftarrow\texttt{env}.\text{step}(a_{t})$
12:  Store $(s_{t},a_{t},r_{t},t)$ in $\mathcal{D}$
13:  if $|\mathcal{D}|\geq W$ and $(t+1)\bmod I=0$
14:    $\hat{Q}=\texttt{FQI}.\text{fit}(\mathcal{D})$
15:  end if
16:end for
Soft Actor-Critic
Soft Actor-Critic (SAC) [haarnoja2018soft, haarnoja2018softa] is an off-policy, model-free algorithm developed to perform entropy-regularized reinforcement learning, where the agent aims to maximize both cumulative return and the entropy of the policy;
$J(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(r_{t}+\alpha H\left(\pi(\cdot\mid s_{t}\right)\right)\right]$
Here, $\alpha$ is the temperature parameter, $H(\cdot)$ is the entropy, and $\gamma$ is the discount factor. Maximizing the entropy of the policy alongside the cumulative reward encourages exploration and prevents premature convergence to suboptimal policies. Algorithm˜3 illustrates the online-SAC algorithm used in this work. We build the time-aware SAC by making the policy, critics, and target critics functions of time $t$.
Algorithm 3 Online Soft Actor-Critic
1:Initialize Foraging environment env, experience buffer $\mathcal{D}$, Policy network $\pi_{\theta}$, Q-function/critic networks $Q_{\phi_{1}}$ and $Q_{\phi_{2}}$, batch size $B$
2:Set target critic parameters equal to main critic parameters $\phi_{\text{targ},1}\leftarrow\phi_{1}$, $\phi_{\text{targ},2}\leftarrow\phi_{2}$
3:for $t=1$ to $T$
4:  $s_{t}\leftarrow\texttt{env}.\text{get\_current\_state()}$
5:  $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t},t)$
6:  $(s_{t+1},r_{t})\leftarrow\texttt{env}.\text{step}(a_{t})$
7:  Store $(s_{t},a_{t},r_{t},t)$ in $\mathcal{D}$
8:  if $|\mathcal{D}|\geq B$
9:    Randomly sample a batch $\mathcal{B}=\{(s_{t},a_{t},r_{t},t)\}$ from $\mathcal{D}$
10:    Compute targets for the critic networks:
$y(r_{t},s_{t+1})=r_{t}+\gamma\left(\min_{i=1,2}Q_{\phi_{\text{targ},i}}(s_{t+1},a^{\prime},t+1)-\alpha\log\pi_{\theta}(a^{\prime}\mid s_{t+1},t+1)\right);\ a^{\prime}\sim\pi_{\theta}(\cdot\mid s_{t+1},t+1)$
11:    Update the critic networks by one step of gradient descent using
$\nabla_{\phi_{\text{targ},i}}\frac{1}{B}\sum_{(s_{t},a_{t},r_{t},t)\in\mathcal{B}}\left(Q_{\phi_{\text{targ},i}}(s_{t},a_{t},t)-y(r_{t},s_{t+1})\right)^{2};\ \text{for}\ i=1,2$
12:    Update policy network by one step of gradient ascent using
$\nabla_{\theta}\frac{1}{B}\sum_{s_{t}\in\mathcal{B}}\left(\min_{i=1,2}Q_{\phi_{\text{targ},i}}(s_{t},a_{\theta}(s_{t}),t)-\alpha\log\pi_{\theta}(a_{\theta}(s_{t})\mid s_{t},t)\right);\ a_{\theta}(s_{t})\sim\pi_{\theta}(\cdot\mid s_{t},t)$
13:    Update the target networks with
$\phi_{\text{targ},i}\leftarrow\rho\phi_{\text{targ},i}+(1-\rho)\phi_{i};\ \text{for}\ i=1,2$
14:  end if
15:end for